OpenStudy (anonymous):

Find the perimeter of a rectangular object which has a length of a square root of 128ft and a width of square root 200ft.

OpenStudy (anonymous):

i got an product of 160 is this correct i think it is but not sure

OpenStudy (anonymous):

45.24?

OpenStudy (anonymous):

square root of 128 is 22.62

OpenStudy (anonymous):

perimeter is length times width so im guessing sqrt 128 times sqrt 200 will give me the correct answer.

OpenStudy (anonymous):

after i't been multiplied by 2

OpenStudy (anonymous):

50.90

OpenStudy (anonymous):

sqrt 128 times sqrt 200 = sqrt 160 then the simplified radical form is 4 sqrt 10 thats wat i got wat do u think?

OpenStudy (anonymous):

\[\sqrt{128} \times \sqrt{200}\]

OpenStudy (anonymous):

\[\sqrt{128} \times \sqrt{200}=\sqrt{160}=4\sqrt{10}\]

OpenStudy (anonymous):

correct? i know how to do it its just i need to verify if this is right cuz the question is kinda long it has 4 parts